Prove that cosec53.cosec37.sec48.sec42 is equal to one ?

Grade:10

1 Answers

Aditya Gupta
5 years ago
sec48=1/cos48=1/sin42 
similarly  cosec53=1/sin53=1/cos37
so, cosec53.cosec37.sec48.sec42=1/[cos37sin37cos42sin42]
=4/sin74sin84, which is not equal  to 1. so question is wrong